A 1.00 g sample of potassium bicarbonate is decomposed by heating. If the resulting potassium carbonate weighs 0.710 g and the calculated yield is 0.690 g, what is the percent yield?
1
Expert's answer
2020-12-01T09:08:30-0500

2KHCO3 = K2CO3 + CO2 + H2O

0.69/138 = 0.005.

0.005 x 100 x 2 = 1.

1-0.69 = 0.31.

1-0.71 = 0.29.

0.29 x 138/62 = 0,64548387 g of K2CO3.

0,64548387 x 100%/0.69= 93,548387%.